OpenStudy (anonymous):

find the power of the complex number (3+i)^4

OpenStudy (anonymous):

4

OpenStudy (anonymous):

it would be 3^4 + i^4

OpenStudy (anonymous):

Use DeMoivre's theorem. First convert \(3+i\) to polar form.

OpenStudy (anonymous):

Yes its 4

OpenStudy (anonymous):

You end up with 81 + i^4

OpenStudy (anonymous):

@mkgarnett Then I suppose \((1+1)^4=2\), because \((1+1)^4=1^4+1^4\) ?

OpenStudy (anonymous):

well i^2 equals 1 @SithsAndGiggles

OpenStudy (anonymous):

my bad it equals -1

OpenStudy (anonymous):

I'm not arguing over what the value of \(i^2\) is. I'm pointing out your mistaken reasoning. \[(a+b)^n\not=a^n+b^n\]

OpenStudy (anonymous):

thats how you do it though.. (a+b)\[\left( a +b \right)^{2} \] equals \[a ^{2} + b ^{2}\]

OpenStudy (anonymous):

or you could even do ab^2

OpenStudy (anonymous):

That's simply not true. \[(a+b)^2=a^2+2ab+b^2\] What you say is true only when one or both of \(a,b\) are 0.

OpenStudy (anonymous):

i thought it was if they dont equal 0

OpenStudy (anonymous):

youre using the foil method

OpenStudy (anonymous):

@Emerson85 convert the given complex number of the form \(x+iy\) into the polar form, \(\large re^{i\theta}\) using \[\begin{cases} x^2+y^2=r^2\\ \tan\theta=\dfrac{y}{x} \end{cases}\] One you have that, use DeMoivre's theorem to raise to the fourth power: \[\large\left(re^{i\theta}\right)^4=r^4e^{4i\theta}=r^4\bigg(\cos4\theta+i\sin4\theta\bigg)\]